Find the length of the hypotenuse of a triangle that has a base lenght of 3, an a side length of 8
Answer by stanbon(75887)

Find the length of the hypotenuse of a triangle that has a base length of 3, an a side length of 8
---
hyp^2 = 3^2 + 8^2
hyp^2 = 9+64 = 73
----
hypotenuse = sqrt(73)
=========================
Cheers,
Stan H.
===========